# references.bib

### v1

@@ 2.3.   Transformer Encoder
The model is a multi-layer transformer-based architecture that serves dual purposes: weakly supervised patient classification and representation learning for unsupervised clustering. It processes sequences of medical concepts derived from patient EHRs and models complex dependencies using stacked transformer encoder layers. Through this architecture, the model produces a probability score for phenotype prediction and simultaneously generates a low-dimensional patient embedding that can be used for downstream tasks such as clustering, subtyping, or visualization.
2.3.1 Multi-Head Attention
Each transformer encoder layer applies multi-head self-attention to capture interactions between medical concepts. Given input embeddings $\mathbf{z}_{k}\in\mathbf{Z}_{i}$, the model computes queries, keys, and values using learnable projection matrices $\mathbf{W}_{Q},\mathbf{W}_{K},\mathbf{W}_{V}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{model}}}$. The self-attention mechanism is split into $H$ heads, each operating in a subspace of dimension $d_{h}=\frac{d_{\text{model}}}{H}$. For each head $h\in\{1,...,H\}$, input embeddings are projected as:
$\mathbf{q}_{k}^{(h)}=\mathbf{W}_{Q}^{(h)}\mathbf{z}_{k},\quad\mathbf{k}_{k}^{(h)}=\mathbf{W}_{K}^{(h)}\mathbf{z}_{k},\quad\mathbf{v}_{k}^{(h)}=\mathbf{W}_{V}^{(h)}\mathbf{e}_{k}^{\text{proj}},$ (6)
$\text{score}_{kj}^{(h)}=\frac{\mathbf{q}_{k}^{(h)}\cdot\mathbf{k}_{j}^{(h)}}{\sqrt{d_{h}}},$ (7)
followed by normalization with the softmax function:
$\alpha_{kj}^{(h)}=\text{softmax}(\text{score}_{kj}^{(h)}),$ (8)
$\mathbf{z}_{k}^{(h)}=\sum_{j}\alpha_{kj}^{(h)}\mathbf{v}_{j}^{(h)}.$ (9)
2.3.2 Concatenation and Linear Projection
$\mathbf{z}_{k}^{\text{attn}}=\text{Concat}(\mathbf{z}_{k}^{(1)},\dots,\mathbf{z}_{k}^{(H)}),\quad\mathbf{z}_{k}^{\text{attn}}\in\mathbb{R}^{d_{\text{model}}}.$ (10)
A final linear projection aggregates information across heads:
$\mathbf{z}_{k}^{\text{out}}=\mathbf{W}^{\text{out}}\mathbf{z}_{k}^{\text{attn}},$ (11)
$\mathbf{z}_{k}^{\text{norm}}=\text{LayerNorm}(\mathbf{z}_{k}+\text{Dropout}(\mathbf{z}_{k}^{\text{out}})).$ (12)
2.3.3 Feedforward Network with SwiGLU Activation
$\mathbf{z}_{k}^{\text{ffn}}=\mathbf{W}^{\text{ffn}}_{2}\text{SwiGLU}(\mathbf{W}^{\text{ffn}}_{1}\mathbf{z}_{k}^{\text{norm}}+\mathbf{b}^{\text{ffn}}_{1})+\mathbf{b}^{\text{ffn}}_{2}.$ (13)
$\mathbf{z}^{*}_{k}=\text{LayerNorm}(\mathbf{z}^{\text{norm}}_{k}+\text{Dropout}(\mathbf{z}_{k}^{\text{ffn}})).$ (14)
@@ 2.4.   Feature Pooling and Fine Tuning
where $\mathbf{W}^{\text{class}}$ and $\mathbf{b}^{\text{class}}$ are learnable parameters. The sigmoid function $\sigma(\cdot)$ maps the logit to a probability in the range $[0,1]$. Model training uses binary cross-entropy (BCE) loss:
$\mathcal{L}_{\text{BCE}}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log(p(y_{i}))+(1-y_{i})\log(1-p(y_{i}))\right],$ (17)
which penalizes confident but incorrect predictions and encourages well-calibrated outputs. After each training round, the best-performing model on the validation set is used to update silver-standard labels using its predicted probabilities:
@@ 3.   Results
We evaluated our framework on two rare pulmonary diseases: PH and severe asthma. For each disease, the model was trained and validated independently using disease-specific cohorts and labels curated from EHR data at BCH.
3.1.   Data Curation
For both PH and severe asthma, we constructed disease-specific cohorts by first identifying an at-risk patient population. The at-risk PH cohort was defined as patients with PheCode 415.2 (suggestive of potential PH), while the severe asthma cohort included patients with PheCodes beginning with J45 (suggestive of asthma of any severity).
The gold-standard cohorts consisted of patients with a confirmed diagnosis, either through expert chart review or enrollment in a disease-specific registry. These high-fidelity labels ($y_{i}^{\text{gold}}$) were used for model evaluation and to calibrate the probabilistic silver-standard labels. The PH gold-standard cohort included 531 patients, and the severe asthma cohort included 248 patients. Each cohort was randomly split into 80% training and 20% test sets, with the test set further divided into two cross-validation folds.
The silver-standard cohorts consisted of the remaining at-risk patients whose phenotype status was not definitively adjudicated – 13,774 for PH and 7,387 for severe asthma. To assign initial probabilistic labels ($y_{i}^{\text{silver}}\in[0,1]$) to these patients, we applied the KOMAP algorithm, an unsupervised phenotyping approach that generates phenotype likelihood scores from EHR data [komap].
For PH, KOMAP was applied to structured, codified EHR data such as diagnosis and procedure codes. For severe asthma, KOMAP incorporated both codified and unstructured clinical data. Natural language features were extracted from clinical notes using two natural language processing (NLP) tools: NILE for named entity recognition [yu2013nile], and GENIE for relation-aware information extraction [ying2025genie]. We specifically targeted mentions of PheCode:415.2 for PH and CUI:C0581126 for severe asthma, corresponding to “severe asthma” in free-text notes.
To improve the interpretability and calibration of KOMAP’s unsupervised outputs, we trained a logistic regression model using the gold-standard subset. This calibration procedure involved three steps: (1) subsetting to patients with gold-standard labels, (2) fitting a logistic model to predict $y_{i}^{\text{gold}}$ from the initial KOMAP scores, and (3) applying the fitted model to the full cohort to generate calibrated probabilities. These values were used as the initial silver-standard labels in our weakly supervised learning framework.
For representation learning, we mapped the codified EHR data in the PH cohort to pre-trained MUGS embeddings [li2024multi] and the NLP-derived features in the severe asthma cohort to pre-trained ONCE embeddings [komap].
@@ 3.2.1 Classification Performance
We evaluated the transformer model using two cross-validation folds, each involving a split of the same 106 gold-standard patients into validation and test subsets. The validation set was used to select the best-performing model checkpoint within each fold, which was then evaluated on an unseen test subset of 53 patients from BCH. Performance metrics – including area under the curve (AUC) and positive predictive value (PPV) – were computed for each fold and averaged. We compared the transformer’s performance to three baselines: (1) a simple rule-based approach based on counts of PH-related PheCodes, (2) the initial silver-standard probabilities produced by KOMAP, (3) a supervised XGBoost model trained on the same features. As shown in Table 1, the transformer outperformed all baselines across key metrics. With sensitivity fixed at 85% for all methods, the transformer achieved the highest AUC (0.913) and PPV (0.876), indicating stronger discriminative performance and more reliable identification of true positives. These results highlight the transformer’s consistent advantages over both knowledge-driven and machine learning-based baselines for phenotyping patients with PH.
Metric Count KOMAP XGBoost Transformer
AUC 0.8450 0.858 0.820 0.913
PPV 0.651 0.808 0.855 0.876
Table 1: Phenotype classification performance for pulmonary hypertension. Transformer metrics are averaged across two cross-validation folds.
@@ 3.2.2 Clustering Performance
To evaluate whether our learned patient representations capture meaningful clinical heterogeneity, we applied t-distributed stochastic neighbor embedding (t-SNE) to reduce the dimensionality of the embeddings and visualize patient distributions [van2008visualizing]. As shown in Figure 2, PH+ and PH- patients are more distinctly separated in the latent space when using transformer-derived embeddings compared to traditional feature engineering approaches. In particular, we compared against embeddings generated using term frequency-inverse document frequency (TF-IDF), a classical technique that weights clinical codes based on their relative frequency across patients [ramos2003using]. While TF-IDF captures code importance in a sparse and interpretable manner, it lacks the contextual awareness and representation capacity of transformer-based models. The improved phenotype separation in the transformer space suggests that our model learns more clinically meaningful patient representations.
Figure 2: t-SNE visualization of patient-level embeddings shows improved separation of PH+ and PH- phenotypes using transformer-based representations.
To further investigate heterogeneity within the PH+ population, we applied principal component analysis (PCA) to the transformer-derived patient embeddings and retained the top components that collectively explained at least 99% of the variance. We then performed $k$-means clustering on the reduced embeddings, identifying two distinct subphenotypes. The first group, referred to as the Slow Progression Cluster ($n=53$), consisted of patients with milder disease trajectories and lower long-term mortality. The second group, labeled the Fast Progression Cluster ($n=16$), included patients with more severe clinical courses and elevated mortality risk. Kaplan–Meier survival analysis revealed a significant difference in 5-year mortality between the two clusters, with the fast progression cluster experiencing markedly worse survival outcomes (log-rank $p=0.0014$; Figure 3).
@@ 3.3.1 Clustering Performance
To investigate heterogeneity among patients with severe asthma, we applied PCA to the transformer-derived patient embeddings, following a similar approach to the PH analysis. We retained the top components explaining at least 99% of the variance and then applied constrained k-means clustering to the reduced embeddings to identify distinct subphenotypes. We subsequently evaluated the resulting clusters with respect to two high-severity outcomes: recurrent low oxygen episodes (e.g., hypoxia, hypoxemia, oxygen desaturation) and respiratory failure. This analysis revealed two clinically meaningful subgroups: a Low-Risk Exacerbator Cluster and a High-Risk Exacerbator Cluster. Patients in the high-risk cluster exhibited a 15.1-fold increased hazard of low oxygen events (hazard ratio [HR] = 15.11, 95% confidence interval [CI]: 2.74–83.31, p = 0.0018) and a 21.6-fold increased hazard of respiratory failure (HR = 21.56, 95% CI: 2.92–159.2, p = 0.0026) compared to those in the low-risk cluster. These findings suggest that the high-risk cluster represents a distinct and more severe asthma trajectory, characterized by frequent respiratory compromise and increased need for acute interventions. This subphenotype may benefit from closer clinical monitoring and more proactive management strategies.
@@ 4.   Conclusion
This work presents a weakly supervised transformer framework for rare disease phenotyping that addresses two central challenges in this research area: limited access to high-quality labels and the need for robust patient representations in heterogeneous EHR data. By combining a small set of gold-standard labels with a large, diverse cohort annotated through probabilistic silver-standard labels, our approach enables effective learning in real-world, label-scarce settings. We demonstrate that the model not only outperforms existing baselines in phenotype classification, but also learns patient embeddings that support meaningful subphenotype discovery and clinical outcome prediction.
Our method contributes a flexible and generalizable strategy for automated phenotyping, integrating EHR-derived embeddings, dynamic label refinement, and transformer-based representation learning in a unified framework. Through case studies in PH and severe asthma, we demonstrate that the transformer model captures nuanced clinical patterns, stratifies disease progression risk, and facilitates downstream tasks such as survival analysis and clustering.
Future directions include enhancing the interpretability of learned representations to support clinical decision-making, validating model performance across multiple institutions and disease areas, and exploring active learning strategies to efficiently incorporate clinician feedback. By bridging algorithmic innovation with practical clinical utility, our framework moves a step closer to enabling scalable, data-driven diagnosis and risk stratification in rare disease populations.
[…]

### v2

@@ 2.3.  Transformer Encoder
Our model builds on a multi-layer transformer encoder but adapts it for the challenges of weakly supervised phenotyping. The encoder serves two purposes simultaneously: (1) patient-level classification, where the model predicts the probability that a patient has the target condition, and (2) representation learning, where it generates low-dimensional embeddings useful for clustering, subphenotyping, and visualization.
Each patient sequence $\mathbf{Z}_{i}$ is processed through stacked transformer encoder layers. Within each layer, multi-head self-attention models dependencies among medical concepts, enabling the network to focus on the parts of the record most informative for the target disease. Standard architectural elements – including residual connections, layer normalization, and feedforward networks with nonlinear activations – are incorporated to ensure stable training. Full mathematical details are provided in Section S1 of the supplementary materials, which describe the multi-layer transformer architecture employed by WEST. The derivations clarify the inner workings of the transformer encoder, including its attention mechanism, projection layers, and feedforward components.
@@ 2.4.  Feature Pooling and Fine Tuning
where $\mathbf{W}^{\text{class}}$ and $\mathbf{b}^{\text{class}}$ are learnable parameters. The sigmoid function $\sigma(\cdot)$ maps the logit to a probability in the range $(0,1)$. Model training employs binary cross-entropy (BCE) loss for classification. After each training round, the best-performing model on the validation set is used to update silver-standard labels using its predicted probabilities:
@@
2.5.  Real World Validation
We evaluated our WEST framework on phenotyping two rare pulmonary diseases, PH and severe asthma, using EHR data from Boston Children’s Hospital. For each disease, the model was trained and validated independently using disease-specific EHR cohorts and labels curated by domain experts via chart review.
2.5.1 Data Curation
For both PH and severe asthma, we constructed disease-specific cohorts by first identifying at-risk patient populations from the EHR data at Boston Children’s Hospital. The at-risk PH cohort comprised 14,305 randomly selected patients with PheCode 415.2 (indicative of potential PH), while the severe asthma cohort comprised 7,822 randomly selected patients with PheCodes beginning with J45 (indicative of asthma of any severity).
Gold-standard cohorts consisted of patients with a confirmed diagnosis, established either through expert chart review or enrollment in a disease-specific registry. The PH gold-standard cohort comprised 531 patients, with 106 (20%) set aside for validation and testing, while the severe asthma cohort comprised 248 patients, with 99 (40%) set aside. These held-out patients were further split into two equally sized cross-validation folds: one used for validation (model checkpoint selection) and the other for testing (performance evaluation). Final performance metrics were averaged across cross-validation folds.
The silver-standard cohorts comprised the remaining at-risk patients whose phenotype status had not been definitively adjudicated – 13,774 for PH and 7,575 for severe asthma. Initial probabilistic labels $y_{i}^{\text{silver}}\in(0,1)$ were assigned to these patients using the KOMAP algorithm [48].
For PH, KOMAP was applied to codified EHR features, including PheCode diagnoses, RxNorm medications, and Clinical Classifications Software (CCS) procedure codes. For severe asthma, KOMAP was applied to natural language features extracted from clinical notes using Narrative Information Linear Extraction (NILE) [53]. From these codes and concepts, we designated PheCode:415.2 for PH and CUI:C0581126 for severe asthma as the target phenotypes. For representation learning, we mapped the codified EHR data in the PH cohort to pre-trained MUGS embeddings [51] and the NLP-derived features in the severe asthma cohort to pre-trained ONCE embeddings [48].
2.5.2 Validation Metrics
We first assessed the classification performance of the WEST pipeline on labels not used during training. Evaluation was performed with two-fold cross-validation, computing AUC, positive predictive value (PPV), and specificity for each fold and then averaging across folds. Sensitivity was fixed at 80%. Performance was compared against five baselines:
  1. 1.
    Count: Frequency of the target concept appearing in each patient’s EHR.
  2. 2.
    KOMAP: The initial silver-standard probabilities generated by KOMAP [48].
  3. 3.
    XGBoost: A supervised gradient-boosted trees classifier [54].
  4. 4.
    Transformer (silver = gold): A transformer trained by treating all silver-standard labels as gold-standard, without any iterative updates or data augmentation.
  5. 5.
    Transformer (gold only): A transformer trained solely on gold-standard labels.
As additional ablation studies, we examined two aspects of gold-standard supervision. First, we varied the number of gold-standard labels used for training, gradually increasing the labeled set from 25 to 400 examples. Second, we modified WEST to train without any gold-standard negative labels, simulating a setting where no confirmed negatives are available and all negative training samples are drawn from the silver-standard cohort.
We next evaluated whether the learned patient representations captured clinically meaningful heterogeneity. Using patients with known disease status who were not included in training, we tested whether the embeddings could distinguish true positive from true negative cases. To visualize this separation, we applied t-distributed stochastic neighbor embedding (t-SNE) to the WEST embeddings of held-out patients and compared the resulting visualization with that obtained from embeddings generated using term frequency-inverse document frequency (TF-IDF), a commonly used feature engineering approach [55, 56].
For subphenotype discovery, we focused on patients predicted to have a positive disease status. We first reduced the learned embeddings using principal component analysis (PCA), retaining components that together explained at least 90% of the variance. We then applied k-means clustering to identify clinically meaningful patient subgroups. Finally, we evaluated the prognostic relevance of these clusters: in PH, by comparing survival distributions using Kaplan-Meier curves; and in severe asthma, by estimating hazard ratios (HRs) for signs and symptoms indicative of disease severity across clusters.
2.6.  Hyperparameter Tuning
We performed hyperparameter tuning using a 2-fold cross-validation procedure to robustly select model configurations. For each hyperparameter setting, the model was trained on one fold and evaluated on the other. A random search strategy was employed to explore the following hyperparameter space: batch size $\in\{64,128,256\}$, learning rate $\in\{5\text{e-}4,1\text{e-}3,2\text{e-}3\}$, hidden dimension $\in\{32,64,128\}$, number of Transformer layers $\in\{2,3,4\}$, dropout rate $\in\{0.3,0.7\}$, and number of training epochs $\in\{30,50\}$. The area under the receiver operating characteristic curve (AUC) served as the primary selection metric, and the chosen hyperparameters for each fold were subsequently used to train the final models.
@@ 3.1.1 Classification Performance
As shown in Table 1, the WEST pipeline trained with both positive and negative gold-standard labels achieved the highest AUC (0.93), PPV (0.95), and specificity (0.92). WEST trained without gold-standard negative labels also outperformed all baselines.
Metric Count KOMAP XGBoost Transformer (silver = gold) Transformer (gold only) WEST (w/o neg) WEST (w/ neg)
AUC 0.85 0.86 0.82 0.84 0.88 0.91 0.93
PPV 0.65 0.88 0.87 0.84 0.89 0.91 0.95
Specificity 0 0.78 0.76 0.70 0.81 0.84 0.92
Table 1: Phenotype classification performance for PH. Transformer and WEST metrics are averaged across two cross-validation folds.
Figures 2 and 7 demonstrate that WEST performance increased with the number of gold-standard training labels. Notably, with only 100 labels, WEST already matched or surpassed all baseline methods, with further gains achieved at larger label sets.
Figure 2: WEST performance as a function of the number of gold-standard training labels. Curves report AUC and F1 score. The horizontal dotted line represents the best performing baseline, Transformer (gold only).
@@ 3.1.2 Clustering Performance
As shown in Figure 5, PH-positive and PH-negative patients were more distinctly separated in the latent space using WEST embeddings than with TF-IDF.
Figure 3: t-SNE visualization of patient-level embeddings for PH phenotypes using TF-IDF and WEST-derived representations.
The WEST pipeline identified 1977 patients with PH. Clustering of PH-positive patient embeddings revealed two subgroups: a Slow Progression cluster ($n=1099$) and a Fast Progression cluster ($n=878$). Kaplan-Meier survival analysis showed a significant difference in 5-year mortality between the two clusters ($p=0.013$; Figure 4).
@@ 3.2.  Severe Asthma
3.2.1 Classification Performance
Table 2 presents classification metrics across methods. When trained with both positive and negative gold-standard labels, WEST achieved the highest AUC (0.87), PPV (0.80), and specificity (0.76) compared against all baselines.
Patients predicted to have severe asthma showed increased risk for multiple indicators of disease severity compared with those predicted to have non-severe asthma. Significant associations were observed for recurrent status asthmaticus (HR = 55.30, 95% CI: 43.93-69.61, $p<0.0001$), respiratory failure (HR = 3.19, 95% CI: 2.05-4.97, $p<0.0001$), low oxygen events including hypoxia and desaturation (HR = 2.66, 95% CI: 2.05-3.45, $p<0.0001$), tachypnea (HR = 3.67, 95% CI: 3.17-4.26, $p<0.0001$), bronchospasm (HR = 3.49, 95% CI: 2.20-5.55, $p<0.0001$), and dyspnea (HR = 2.97, 95% CI: 2.66-3.32, $p<0.0001$) (Figure 6).
Metric Count KOMAP XGBoost Transformer (silver = gold) Transformer (gold only) WEST (w/o neg) WEST (w/ neg)
AUC 0.80 0.82 0.83 0.82 0.78 0.86 0.87
PPV 0.74 0.74 0.78 0.70 0.69 0.75 0.80
Specificity 0.68 0.66 0.72 0.60 0.60 0.69 0.76
Table 2: Phenotype classification performance for severe asthma. Transformer metrics are averaged across two cross-validation folds.
@@ 3.2.2 Clustering Performance
Again, patients with and without severe asthma were more distinctly separated in the latent space when using WEST embeddings compared to TF-IDF embeddings (Figure 5).
Figure 5: t-SNE visualization of patient-level embeddings for severe asthma (SA) phenotypes using TF-IDF and WEST-derived representations.
Among 582 patients predicted to have severe asthma, k-means clustering identified a Low Exacerbator cluster ($n$ = 209) and a High Exacerbator cluster ($n$ = 373). Patients in the High Exacerbator cluster had higher risk of recurrent status asthmaticus (HR = 2.35, 95% CI: 1.91-2.91, $p<0.0001$), respiratory failure (HR = 2.68, 95% CI: 1.31-5.47, $p=0.0068$), low oxygen events (HR = 1.54, 95% CI: 1.05-2.28, $p=0.0291$), and tachypnea (HR = 1.41, 95% CI: 1.11-1.79, $p=0.0050$) compared with the Low Exacerbator cluster (Figure 6).
Figure 6: Hazard ratios and 95% CIs for adverse events associated with asthma severity and exacerbation subphenotypes.
4.  Discussion
In this study, we introduce WEST, a weakly supervised transformer framework that integrates gold-standard annotations with iteratively refined silver-standard labels to enable data-efficient rare disease phenotyping from heterogeneous EHR data. Across PH and severe asthma case studies, WEST consistently outperformed rule-based and conventional machine learning baselines, producing patient representations that more clearly separated disease states and revealed clinically meaningful subgroups with divergent outcomes.
Among PH patients, clustering of WEST embeddings uncovered two distinct subgroups that we termed the Slow Progression and Fast Progression clusters. Patients in the latter group exhibited significantly poorer survival, highlighting clinically relevant heterogeneity in disease trajectories that may not be captured by current diagnostic labels. In severe asthma, clustering differentiated Low Exacerbator and High Exacerbator subgroups, the latter showing increased risks of recurrent status asthmaticus, respiratory failure, hypoxemia, and tachypnea. These patterns suggest that WEST-derived representations encode latent disease biology and care trajectories that extend beyond codified diagnoses. Such properties position WEST as a potential foundation for precision registry curation, proactive monitoring, and targeted intervention, enabling earlier identification of high-risk subgroups and more efficient allocation of clinical resources.
Methodologically, our work contributes to the growing literature on weak supervision and transformer-based modeling in healthcare, demonstrating that iterative silver-standard label refinement can effectively reduce noise and improve calibration during training. This label-efficient paradigm balances the breadth of routinely collected EHR data with the reliability of expert-validated annotations, yielding performance gains over both silver-only and fully supervised baselines. The superiority of transformer embeddings over TF-IDF representations further underscores the value of contextual modeling: by capturing temporal and semantic dependencies among codes, WEST generates clinically coherent patient spaces that facilitate downstream discovery, risk stratification, and hypothesis generation.
Several limitations merit discussion. First, our evaluation was retrospective and based on data from a single health system, which may limit generalizability to other care settings or populations. Second, while WEST integrates structured and unstructured data, the current implementation focuses on single-phenotype classification. Future extensions to multi-task learning and cross-institutional training could enhance scalability and robustness, particularly for ultra-rare diseases. Lastly, prospective validation and clinician-in-the-loop evaluation will be essential to assess real-world utility and interpretability in clinical workflows.
In summary, WEST advances weakly supervised learning for rare disease phenotyping by coupling transformer-based representation learning with iterative silver-label refinement. This framework improves data efficiency, enhances phenotype separability, and captures latent clinical heterogeneity, offering a scalable approach to harness EHR data for precision discovery and translational impact.
@@ 5.  Conclusion
This study demonstrates that integrating weak supervision with transformer-based modeling enables scalable and data-efficient phenotyping for rare diseases. By coupling limited expert annotations with iteratively refined supervision, the framework achieves robust performance while capturing clinically meaningful heterogeneity within PH and severe asthma cohorts. These findings illustrate how weakly supervised transformers can complement traditional EHR-based phenotyping approaches by leveraging routine clinical data to reveal structure that extends beyond diagnostic codes and by supporting more consistent identification of patients with underrecognized or misclassified rare diseases.
More broadly, this work illustrates a general strategy for studying conditions that are too rare or heterogeneous for conventional supervised learning. Rather than replacing expert review, weakly supervised transformers can serve as augmentation tools – helping prioritize cases for validation, enrich registries, and identify underrecognized subgroups for further study. Embedding such approaches within data curation and clinical research pipelines can accelerate discovery, enhance diagnostic accuracy, and improve the representativeness of real-world evidence for rare disease populations.
6.  Reproducibility
6.1.  Data Availability
The EHR data used in this study were obtained from Boston Children’s Hospital and contain protected health information that cannot be shared publicly due to patient privacy regulations and institutional data use agreements. Access to these data is therefore restricted and cannot be distributed outside the institution. Derived, de-identified summary results supporting the findings of this study are available from the corresponding author upon reasonable request and subject to institutional approval.
6.2.  Code Availability
Python implementation of the methodology developed and used in this study is available on GitHub at https://github.com/kfgreco/WEST.
Supplementary Materials
S1.  Mathematical Details for Transformer Encoder
Here, we provide additional mathematical details on the multi-layer transformer architecture employed by WEST. These derivations clarify the inner workings of the transformer encoder, including its attention mechanism, projection layers, and feedforward components.
S1.1.  Multi-Head Attention
Each transformer encoder layer applies multi-head self-attention to capture interactions between medical concepts. Given input embeddings $\mathbf{z}_{ki}\in\mathbf{Z}_{i}$, the model computes queries, keys, and values using learnable projection matrices $\mathbf{W}_{Q},\mathbf{W}_{K},\mathbf{W}_{V}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{model}}}$. The self-attention mechanism is split into $H$ heads, each operating in a subspace of dimension $d_{h}=\frac{d_{\text{model}}}{H}$. For each head $h\in\{1,...,H\}$, input embeddings are projected as:
$\mathbf{q}_{ki}^{(h)}=\mathbf{W}_{Q}^{(h)}\mathbf{z}_{ki},\qquad\mathbf{k}_{ki}^{(h)}=\mathbf{W}_{K}^{(h)}\mathbf{z}_{ki},\qquad\mathbf{v}_{k}^{(h)}=\mathbf{W}_{V}^{(h)}\mathbf{e}_{k}^{\text{proj}},$ (S1.1)
$\text{score}_{kji}^{(h)}=\frac{\mathbf{q}_{ki}^{(h)}\cdot\mathbf{k}_{ji}^{(h)}}{\sqrt{d_{h}}},$ (S1.2)
followed by normalization with the softmax function:
$\alpha_{kji}^{(h)}=\text{softmax}(\text{score}_{kji}^{(h)}),$ (S1.3)
$\mathbf{z}_{ki}^{(h)}=\sum_{j}\alpha_{kji}^{(h)}\mathbf{v}_{j}^{(h)}.$ (S1.4)
S1.2.  Concatenation and Linear Projection
$\mathbf{z}_{ki}^{\text{attn}}=\text{Concat}(\mathbf{z}_{ki}^{(1)},\dots,\mathbf{z}_{ki}^{(H)}),\qquad\mathbf{z}_{ki}^{\text{attn}}\in\mathbb{R}^{d_{\text{model}}}.$ (S1.5)
A final linear projection aggregates information across heads:
$\mathbf{z}_{ki}^{\text{out}}=\mathbf{W}^{\text{out}}\mathbf{z}_{ki}^{\text{attn}},$ (S1.6)
$\mathbf{z}_{ki}^{\text{norm}}=\text{LayerNorm}(\mathbf{z}_{ki}+\text{Dropout}(\mathbf{z}_{ki}^{\text{out}})).$ (S1.7)
S1.3.  Feedforward Network with SwiGLU Activation
$\mathbf{z}_{ki}^{\text{ffn}}=\mathbf{W}^{\text{ffn}}_{2}\text{SwiGLU}(\mathbf{W}^{\text{ffn}}_{1}\mathbf{z}_{ki}^{\text{norm}}+\mathbf{b}^{\text{ffn}}_{1})+\mathbf{b}^{\text{ffn}}_{2}.$ (S1.8)
$\mathbf{z}^{*}_{ki}=\text{LayerNorm}(\mathbf{z}^{\text{norm}}_{ki}+\text{Dropout}(\mathbf{z}_{ki}^{\text{ffn}})).$ (S1.9)
S2.  Supplementary Ablation Results (PPV and Specificity)
Figure 7: WEST performance as a function of the number of gold-standard training labels. Curves report PPV and specificity. The horizontal dotted line represents the best performing baseline, Transformer (gold only).